1)For a monopolistic firm demand and total cost (TC) functions are as follows:
q = 230 – (1/2) P and TC = 20 + (1/2) q^2
What is the equilibrium level of P (P*) that maximizes the profit of firm?

2)) If the total cost function (TC) of a competitive firm is TC = 20 + 0.2 Q^2 and the price is (P) = 400 what is the equilibrium level of quantity (Q*) that maximizes profit?
1
Expert's answer
2020-07-29T10:58:47-0400

1) Answer

"P^* = \$276"


Solution

For the monopolistic firm to maximize profit, it will produce at the level of output where "MR = MC."


From the demand function "q = 230 - \dfrac {1} {2} p" , the inverse demand function is found by making "p" the subject of formula.


"=> 2q = 460 - p"

"=> p = 460 - 2q"

But, "P = AR"

"=> AR = 460 - 2q"


"TR = AR \u00d7 q"

"= (460 - 2q) \u00d7 q"

"= 460q - 2q^2"


"MR = \dfrac {d} {dq} (TR)"

"= \dfrac {d}{dq} (460q - 2q^2)"

"= 460 - 4q"


Also,

"MC = \dfrac {d}{dq} (TC)"

"= \dfrac {d}{dq} (20 + \dfrac {1}{2} q^2)"

"= q"


Now,

"MR = MC"

"=> 460 - 4q = q"

"460 = 4q + q"

"460 = 5q"

"\therefore q = 92 \space units"


Substituting 92 units for q in the inverse demand function gives:


"p = 460 - 2(92)"

"= 460 - 184"

"= \$276"


Therefore, the equilibrium level of price that maximizes the firm's profit, "P^* = \$276"


2) Answer

"Q^* = 1,000 \space units"


Solution

The profit maximizing condition is "MR = MC."


For a competitive firm, "P = AR = MR"

"\therefore \space since \space P = \$400,\space it \space follows \space that \space MR = \$400"

Now,

"MC = \dfrac {d}{dQ} (TC)"

"= \dfrac {d} {dQ} (20 + 0.2Q^2)"

"= 0.4Q"


"Since \space MR = MC,"

"=> \$400 = 0.4Q"

"=> Q = 1,000 \space units"


"\therefore" the equilibrium level of output that maximizes the firm's output, "Q^* = 1,000 \space units"